### v1

@@ 2 Related Work
The pursuit of efficient attention mechanisms has been a central focus in deep learning research, driven by the increasing scale of Transformer-based models and their application to long-sequence tasks [2, 3, 4, 5, 7, 6]. This section reviews key developments in attention optimization, emphasizing their relevance to inference on mid-tier GPUs, with particular attention to the challenges posed by hardware like the NVIDIA H20, which is limited by its 148 TFLOPS FP16 compute capability.
@@ 3.2.2 FlashMLA-ETAP: Algorithm and Implementation
11:         Compute $\widetilde{\mathbf{P}}_{i}^{(j)}=\exp(\mathbf{S}_{i}^{(j)}-m_{i})$ and $\ell_{i}=\exp(m_{i}^{\text{old}}-m_{i})\ell_{i}+\text{rowsum}(\widetilde{\mathbf{P}}_{i}^{(j)})$.
@@ 3.2.3 Explanation of FlashMLA-ETAP Pseudocode
This implementation minimizes padding overhead by aligning the KV context length with the $M$-dimension, achieving significant performance improvements on the H20 while maintaining compatibility with FlashMLA’s low-rank compression and inference optimizations.
@@ 3.2.4 Theoretical Analysis of ETAP’s Integration into FlashAttention-3 and FlashInfer
To substantiate the claim that ETAP can be seamlessly integrated into other attention frameworks such as FlashAttention-3 and FlashInfer, we provide a theoretical analysis of its compatibility with their computational pipelines, focusing on the structural similarities and differences that influence integration feasibility. Both FlashAttention-3 and FlashInfer, like FlashMLA, are designed to optimize attention mechanisms on modern GPUs, but they differ in their implementation strategies and target hardware optimizations, necessitating a careful examination of how ETAP’s transposition-based approach aligns with their architectures.

### v2

@@ 2 Related Work
The pursuit of efficient attention mechanisms has been a central focus in deep learning research, driven by the increasing scale of Transformer-based models and their application to long-sequence tasks [2, 3, 4, 15, 16, 5, 7, 6]. This section reviews key developments in attention optimization, emphasizing their relevance to inference on mid-tier GPUs, with particular attention to the challenges posed by hardware like the NVIDIA H20, which is limited by its 148 TFLOPS FP16 compute capability.
@@ 3.2.2 FlashMLA-ETAP: Algorithm and Implementation
11:         Compute $\widetilde{\mathbf{P}}_{i}^{(j)}=\exp(\mathbf{S}_{i}^{(j)}-m_{i})$ and $\ell_{i}=\exp(m_{i}^{\text{old}}-m_{i})\ell_{i}+\text{colsum}(\widetilde{\mathbf{P}}_{i}^{(j)})$.
@@ 3.2.3 Explanation of FlashMLA-ETAP Pseudocode
This implementation minimizes padding overhead by aligning the KV context length with the $M$-dimension, achieving significant performance improvements on the NVIDIA H20 while maintaining compatibility with FlashMLA’s low-rank compression and inference optimizations.
@@ 3.2.4 Theoretical Analysis of ETAP’s Integration into FlashAttention-3 and FlashInfer
To substantiate the claim that ETAP can be seamlessly integrated into other attention frameworks such as FlashAttention-3 [10] and FlashInfer [13], we provide a theoretical analysis of its compatibility with their computational pipelines, focusing on the structural similarities and differences that influence integration feasibility. Both FlashAttention-3 and FlashInfer, like FlashMLA, are designed to optimize attention mechanisms on modern GPUs, but they differ in their implementation strategies and target hardware optimizations, necessitating a careful examination of how ETAP’s transposition-based approach aligns with their architectures.